OpenStudy (anonymous):

The sum of the deviations of a set of n scores \[x_{_{1}},x_{_{2}}..............,x_{_{n}}\] measured from 50 is -10 and when measured from 46 is 70.Then the mean is

OpenStudy (anonymous):

Let \( S= \sum_{i=1}^n x_i \) then \[ S-50 n =-10\\ S- 46 n = 70 \] This will give you S and n and the rest is easy.

OpenStudy (anonymous):

You get n=20 and S =990 So the mean is \[ \frac {990}{20}=\frac {99}2=49.5 \]